How do you solve: x-6sqrtx+8=0
and
(x^2-6x)^2-22(x^2-6x)-135=0

Answer by edjones(8007)

x-6sqrt(x)+8=0
x+8=6sqrt(x)
x^2+16x+64=36x square each side.
x^2-20x+64=0
(x-16)(x-4)=0
x=16, x=4
.
Ed